A gardener charges £25 for each visit and £15 an hour for every hour she works. How much does she charge if she works for 4 hours
Answer by Theo(13342)

the equation is:

r = 25 + 4 * x

r is the revenue
x is the number of hours worked.

if the gardener works 4 hours, then the total revenue is 25 + 4 * 15 = 25 + 60 = 85